# Predicting from Strings: Language Model Embeddings for Bayesian Optimization

### v2

@@ Abstract
Bayesian Optimization is ubiquitous in the field of experimental design and blackbox optimization for improving search efficiency, but has been traditionally restricted to regression models which are only applicable to fixed search spaces and tabular input features. We propose Embed-then-Regress, a paradigm for applying in-context regression over string inputs, through the use of string embedding capabilities of pretrained language models. By expressing all inputs as strings, we are able to perform general-purpose regression for Bayesian Optimization over various domains including synthetic, combinatorial, and hyperparameter optimization, obtaining comparable results to state-of-the-art Gaussian Process-based algorithms. Code can be found at github.com/google-research/optformer/embed_then_regress.
@@ 1 Introduction
A fundamental component of all value-based search methods is regression, in which proposed solutions are filtered by predictions on their performance, before evaluation. By utilizing an accurate regression model, or regressor, along with balanced explore-exploit mechanisms, large improvements to the sample complexity of search have been widely possible. However, many regression methods so-far have been task-specific, due to the reliance of modelling assumptions and dimensionality constraints. Learning-based regression methods are particularly susceptible to this issue, due to their reliance on fixed-length tensors for input representation.
@@
Recent progress in large language models (LLMs) have demonstrated the flexibility and versatility of representation of information as strings, which allow for a wider range of data formats to be encoded for subsequent processing. The potential of LLMs for universal learning-based regression is considerable, allowing for regressors that can be generalized across multiple tasks, thereby mitigating the task-specific limitations of current methods.
In this work we focus on improving the flexibility of regressor-guided search methods through the use of LLM-based embeddings, which map arbitrary strings to fixed-length vectors to be used in downstream tensor-based regressor models, such as an in-context learning (ICL) based Transformer. Specifically, our contributions are:
  • Our framework “embed-then-regress” is one of the first to allow the use of free-form string representations in Bayesian Optimization, by using language models to embed trial inputs as features for in-context regression.
  • By pretraining a Transformer-based regressor at scale over a large variety of offline evaluation data, we can achieve uncertainty-aware numeric predictions over unseen objective functions.
  • After augmenting the framework with explore-exploit techniques, we achieve competitive optimization results over a variety of optimization tasks, including synthetic, combinatorial, and hyperparameter optimization.
@@ 2 Related Work and Motivation
Bayesian Optimization refers to a class of techniques which use regressors for solving blackbox optimization problems, by suggesting best candidates according to an explore-exploit tradeoff. For a given objective, the speed of optimization relies heavily on the regressor’s underlying prior, or assumptions about the nature of the objective, such as its smoothness and landscape. Largely dominated by the use of Gaussian Process (GP) regressors, the field of Bayesian Optimization has thus seen a rise in works (Wang et al., 2024; Fan et al., 2024) which seek to learn better prior GP hyperparameters such as length-scales and kernel amplitudes based on offline pretraining or manually designed feature representations for combinatorial objects (Deshwal et al., 2023), while keeping underlying kernel definitions fixed.
@@
Numerous end-to-end neural network-based approaches such as the use of attention mechanisms and Transformers (Vaswani et al., 2017) have been introduced to allow more learnable behaviors, and we refer the reader to (Song et al., 2024b) which provides a general reference on their use for blackbox optimization. Relevant to our particular case of regression, works such as (Nguyen and Grover, 2022; Garg et al., 2022) demonstrated the benefits of using raw Transformers as in-context regression models, or neural processes, with others (Bai et al., 2023; Zhang et al., 2024) establishing provable guarantees. Similarly, (Müller et al., 2022) demonstrated that Transformers trained on synthetic data as “prior-fitted networks” are capable of Bayesian inference, leading to their use in Bayesian optimization (Müller et al., 2023; Nguyen et al., 2023; Nguyen and Grover, 2024).
@@
Unfortunately, as both raw Transformers and GPs require fixed dimensional features, this limits their applications to inputs expressable as tabular features for e.g. hyperparameter tuning, or task-specific embeddings for e.g. chemistry (Maus et al., 2022). Further works have attempted to improve the flexibility of regression-modeling through the use of token-based representations, which allows regressors to be simultaneously used over various types of input formats. Since the context-window of Transformers still remain the most expensive limitation, a useful organization of recent works can be based on their treatment of the total sequence length, roughly equal to:
[…]
@@
Among sequence-to-sequence methods which pretrain for blackbox optimization specifically, (Chen et al., 2022) uses custom tokenizations to minimize trial token length in order to maximize trial count. However, this is restricted to very constrained search spaces (e.g. flat hyperparameter spaces), and lacks flexibility in utilizing arbitrary forms of data. In contrast, (Song et al., 2024a) allows arbitrary string representations, but the large token length of each trial severely limits the trial count allowed in the context window, forcing the use of alternative but tedious methods of absorbing online data, such as inference-time fine-tuning.
Other methods (Vacareanu et al., 2024) use text-to-text chat-based services such as ChatGPT (OpenAI, 2022) and Gemini (Google, 2024) to demonstrate their emergent capabilities for in-context regression, but such methods lack the ability to pretrain over large amounts of offline evaluations. Efforts in ICL-based reward modeling with chat-based LLMs (Lightman et al., 2024) allow fine-tuning over chains of thought, but have only used coarse discretized scores of e.g. $\{-1,0,1\}$ rather than highly precise numeric predictions over vastly different scales.
@@
  • Production of precise numeric predictions over diverse objective scales.
This naturally leads to the use of embedding-based methods which can compress any string representation of a candidate into a feature vector, using only a single unit of sequence length when sent to a ICL model such as a raw Transformer. This can be seen as a form of encoding and decoding, which has been shown to produce competitive results over pure language modeling tasks (Raffel et al., 2020).
@@ 3.1 Preliminaries
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a real-valued function over a search space $\mathcal{X}$. The goal of blackbox optimization is to produce an $x^{*}$ which maximizes $f$:
@@
We define a regressor as a prediction model which can output a distribution of prediction values for $f(\cdot)$ over a query point $x$, given the history $\{x_{s},y_{s}\}_{s=1}^{t}$ of trajectory of $t$ evaluations over $f$ so far. Such regressors may also be learnable over additional offline data besides the given history.
During inference, the regressor may be turned into an acquisition function $a:\mathcal{X}\rightarrow\mathbb{R}$ to represent explore-exploit tradeoffs. We assume the existence of a (potentially task-dependent) acquisition optimizer which can quickly and cheaply sample suggestions $x\in\mathcal{X}$, usually in the form of an evolutionary algorithm. The history-dependent acquisition $a_{t+1}(\cdot)$ may thus be used to filter out poor samples, or used in an entire Bayesian optimization loop in which the acquisition optimizer is used to find $x_{t+1}:=\operatorname*{arg\,max}_{x\in\mathcal{X}}a_{t+1}(x)$ as the next $x$-proposal.
@@ 3.2 In-context Transformer Regressor
For our underlying regression model, we then use an additional Transformer (Vaswani et al., 2017), by sending in as input sequence $(\overline{x}_{1}\oplus\overline{y}_{1}),\ldots,(\overline{x}_{t}\oplus\overline{y}_{t})$ where $\overline{y}\in\mathbb{R}^{d}$ is the feature representation of the float $y$ after applying a trainable projection, and $\overline{x}\oplus\overline{y}$ is the trial representation expressed as the concatenation of $\overline{x}$ and $\overline{y}$.
@@
In order to obtain a prediction for a query point $x$, we may then further append a query $(\overline{x}\oplus\overline{0})$ to the history input sequence where $\overline{0}$ is a dummy value, and following a forward pass of the Transformer where $(\overline{x}\oplus\overline{0})$ attends to all previous trials, post-process the corresponding $t+1$-th output feature with a parametric output distribution over $\mathbb{R}$. For our case we assume a Gaussian $\mathcal{N}(\mu_{t+1}(x),\sigma^{2}_{t+1}(x))$ using mean and standard deviation output heads. The key components to our method can be summarized in Figure 2.
@@
Parallel Predictions: In order to simultaneously predict over a given set of $k$ target points $x_{t+1},\ldots,x_{t+k}$, we additionally append $(\overline{x}_{t+1}\oplus\overline{0}),\ldots,(\overline{x}_{t+k},\overline{0})$ to the history sequence and generate a custom attention pattern of shape $(t+k)\times t$ where all tokens to attend to the history while no tokens attend to the targets. This allows an efficient parallel modeling of $p(y_{t+i}\>|\>x_{t+i},\{x_{s},y_{s}\}_{s=0}^{t})$ for $1\leq i\leq k$ which speeds up training when computing a summed loss over multiple target predictions.
$y$-Normalization: Depending on the function, $y$-value outputs may contain a wide variety of scales, and need to be normalized properly. We can define a normalization procedure parameterized by a history of objectives $(y_{1},\ldots,y_{t})$, also applicable to incoming target values. These steps consist of, in order: (1) Shifting objectives to have zero mean and divide by standard deviation. (2) Reducing harmful effects of bad outliers by fitting the “bad half” of objectives $\{y_{i}\leq y_{\text{median}}\}$ to a normal curve, using percentiles as z-scores. (3) Linearly scale $y\leftarrow\frac{y-y_{\text{min}}}{y_{\text{max}}-y_{\text{min}}}$ which ensures all historical $y$-values within $[0,1]$ and apply additional damping (e.g. sigmoid or log transform) to target values significantly outside this range.
@@ 3.3 Pretraining and Inference
While the embedder is frozen, we pretrain the weights $\theta$ of the ICL regression Transformer, over all such offline evaluation data. Each training example consists of a sampled task and history cutoff length $t^{\prime}\in[0,T)$ so that $\{x_{s},y_{s}\}_{s\leq t^{\prime}}$ is considered a history, while $\{x_{t^{\prime}+i},y_{t^{\prime}+i}\}_{t^{\prime}+i}^{T}$ are target points, with the loss computed as the sum of prediction losses over all targets, i.e.
$\sum_{i=1}^{T-t^{\prime}}\ell_{\theta}(x_{t^{\prime}+i},y_{t^{\prime}+i};\{x_{s},y_{s}\}_{s=1}^{t^{\prime}})$ (3)
@@ 3.4 Model Details
In this paper, to demonstrate the validity of our approach on relatively low compute budgets, we intentionally use relatively smaller language model embedder sizes in comparison to the larger GPT (OpenAI, 2023) or Gemini (Google, 2024) family of models. Specifically, we use a pretrained T5-XL encoder (1B parameters), based on the encoder-decoder T5-family of models (Raffel et al., 2020). Along with only 8 layers of the underlying regression Transformer, this leads to a maximum required training budget of approximately 16 GPUs for training and 1 GPU for inference, possible with most academic budgets. Further details with respect to model sizes and training hyperparameters can be found in Appendix A.
@@ 4 Experiments
4.1 End-to-End Blackbox Optimization
To emphasize the broad applicability of our Embed-then-Regress method, our emphasis is not on achieving state-of-the-art results compared to domain-specific baselines, but rather demonstrating its general applicability across a variety of tasks. We outline possible improvements within specific domains in Section 5. We evaluate the performance of our algorithm on various problems consisting of synthetic, combinatorial, and hyperparameter optimization (exact details in Appendix B). All curves are plotted with mean and 0.5 deviation as error bars over 10 repeats.
Synthetic Optimization: In common optimization scenarios, the search space is a flat Cartesian product of float and categorical parameter types. Our string-based regression will represent each $x$ with standard JSON over the dictionary mapping parameter names to values, e.g. for a search space with two parameters, one continuous named p0 and another integer p1, the string representation for an example trial would be {"p0":0.3,"p1":4}.
@@ 4.1 End-to-End Blackbox Optimization
We benchmark over the Blackbox Optimization Benchmarking (BBOB) suite (ElHara et al., 2019), one of the most widely used synthetic function benchmarks, containing 24 different objectives over continuous search spaces. In order to have a notion of offline “training” data and unseen “test” functions to be optimized, we split the original functions across each landscape type (Separable, Ill-Conditioned, etc.) into training and test sets, and additionally apply randomized transformations (e.g. shifting, rotating, discretizing, increasing/decreasing dimensions) over all objectives to induce non-continuous search spaces with categorical parameters and avoid overfitting. Evaluations used for offline training data were uniformly sampled over the search space.
As a traditional GP baseline, we use the industry-grade UCB-based Bayesian Optimization method “GP-Bandit” (Song et al., 2024c) from Open Source Vizier (Song et al., 2022). In order to control for confounding factors affecting performance, our method uses the same “Firefly” acquisition optimizer as Vizier’s, with the same evaluation budget. In Figure 3, we find that Embed-then-Regress is generally comparable with and interestingly can even significantly outperform GP-Bandit in a few cases.
@@
Each of these problems can be parameterized by a set of coefficients (e.g. city locations for Travelling Salesman, matrix entries for log-determinant). Note that we are in the bandit setting, in which these coefficients are hidden from the algorithm and the only feedback is the final objective. Similar to before, we thus can also generate offline pretraining data by randomizing these coefficients and problem sizes, and evaluating over random candidates. For our string regression, we may simply use JSON over indices; e.g. {[0]:2, [1]:0, [2]:3, [3]:1} for a permutation space of size 4, e.g. {[0]:1, [1]:3} for a (${4\choose 2}$ choice space.
@@
Figure 4: ($\uparrow$) Higher is better. Best-so-far curves across 8 randomized combinatorial problems. Title parenthesis $(P)$ means a permutation space of size $P$ and $(N,K)$ denotes a (${N\choose K}$ choice space. Note that plotting begins at trial 20, since previous trials are random.
@@
While there are few previous works using GPs for e.g. permutation spaces (Deshwal et al., 2022; Oh et al., 2022), they require constructing very domain-specific kernels and complex acquisition optimizers (e.g. semi-definite programming) making them difficult to reproduce. We thus use a simpler optimizer such as Regularized Evolution (Real et al., 2019) which does not need modelling assumptions other than implementing random mutations between trials and can be used broadly (Real et al., 2020). We also empirically found this was better than other evolutionary alternatives such as NSGA-II (Deb et al., 2002) or hill-climbing.
@@
Rather than fully optimizing the acquisition which be slow for large-scale evolutionary searches, we can simply apply best-of-many sampling by using the regressor’s UCB acquisition to rank sample candidates proposed by evolution, and suggest only the best. In Figure 4, we see that this boosts exploration over the original Regularized Evolution, which can often get stuck at local optima early on.
@@
Figure 5: ($\uparrow$) Higher is better. Best-so-far curves over 8 randomly chosen hyperparameter surrogate functions. Title contains task summary along with number of parameters (#$(\#P)$. Normalized objective values are displayed since raw objective values over large ranges, e.g. $y\in[-10^{7},10^{7}]$ from private functions would lack meaningful context.
Hyperparameter Optimization: With the advent of industry-wide hyperparameter tuning services (Golovin et al., 2017; Balandat et al., 2020; Liaw et al., 2018), large amounts of evaluations over expensive but realistic experiments can be stored offline. To efficiently benchmark our method over objectives encountered in the wild, we use surrogate-based benchmarking which has shown to be effective (Zela et al., 2022; Eggensperger et al., 2015) for providing realistic yet cheap comparisons between different algorithms without using huge amounts of compute. Without surrogates, benchmarking could take weeks in wall-clock time and hundreds to thousands of GPU-hours.
@@
Figure 6: ($\uparrow$) Higher is better. Violin plots of log-efficiency scores across algorithms over 20 hyperparameter optimization (HPO) surrogate objectives. Embed-then-Regress (EtR) as the reference algorithm is displayed horizontally at $x=0$.

### v3

# Language Model Embeddings Can Be Sufficient for Bayesian Optimization

@@ Abstract
Bayesian Optimization is ubiquitous in experimental design and black-box optimization for improving search efficiency. However, most existing approaches rely on regression models which are limited to fixed search spaces and structured, tabular input features. This paper explores the use of LLM embeddings over string inputs for in-context regression in Bayesian Optimization. Our results show that representing inputs as strings enables general-purpose regression across diverse domains, including synthetic, combinatorial, and hyperparameter optimization. Furthermore, our approach achieves optimization performance comparable to state-of-the-art Gaussian Process-based methods such as Google Vizier, and demonstrates potential for broader and more flexible applications.
@@ 1 Introduction
A fundamental component of all value-based search methods is regression, in which proposed solutions are filtered by predictions on their performance, before evaluation. By utilizing an accurate regression model, or regressor, along with balanced explore-exploit mechanisms such those proposed in Bayesian Optimization, large improvements to the sample complexity of search have been widely possible. However, for the field of traditional optimization, many regression methods so-far have been task-specific, due to the reliance of modeling assumptions and dimensionality constraints. Common tabular regression methods such as random forests and Gaussian Processes are particularly susceptible to this issue, as their input dimensions are dependent on the problem space.
@@
Large language models (LLMs) offer a promising approach to regression due to their ability to represent information as strings, enabling more flexible input formats than traditional methods. Unlike tabular models, LLMs are not constrained by fixed-length input representations, making them more adaptable across different tasks. This flexibility allows LLM-based regressors to generalize beyond task-specific settings, addressing one of the key limitations of existing regression techniques. If LLMs prove effective in optimization, they could have far-reaching implications for accelerating search across a wide range of evaluation functions that map strings to objective values.
As a starting point, in this work we focus on the well-established setting of traditional black-box optimization over tabular inputs. We explore the use of LLM-based embeddings, which map arbitrary strings into fixed-length vectors suitable for downstream tensor-based regressors that effectively balance exploration and exploitation. Specifically, our contributions are as follows:
  • Our framework "embed-then-regress" describes a simple recipe for using free-form string representations in Bayesian Optimization, by using language models to embed trial inputs as features for in-context regression.
  • By using a standard Transformer Neural Process (Nguyen and Grover, 2022) as the downstream regressor and pretraining it at scale over a large variety of offline evaluation data, we can achieve accurate and well-calibrated numeric predictions over unseen objective functions, competitive with specialized, state-of-the-art methods.
  • After augmenting the framework with explore-exploit techniques, we achieve competitive optimization results over a variety of optimization tasks, including synthetic, combinatorial, and hyperparameter optimization, in particular matching the performance of industry-grade traditional Gaussian Process techniques such as Google Vizier.
@@ 2 Related Work and Motivation
Bayesian Optimization refers to a class of techniques which use regressors for solving non-differentiable optimization problems, by suggesting best candidates according to an explore-exploit tradeoff. For a given objective, the speed of optimization relies heavily on the regressor’s underlying prior, or assumptions about the nature of the objective, such as its smoothness and landscape. Largely dominated by the use of Gaussian Process (GP) regressors, the field of Bayesian Optimization has thus seen a rise in works (Wang et al., 2024; Fan et al., 2024) which seek to learn better prior GP hyperparameters such as length-scales and kernel amplitudes based on offline pretraining or manually designed feature representations for combinatorial objects (Deshwal et al., 2023; White et al., 2021; Ru et al., 2021), while keeping underlying kernel definitions fixed.
@@
Numerous end-to-end neural network-based approaches such as the use of attention mechanisms and Transformers (Vaswani et al., 2017) have been introduced to allow more learnable behaviors, and we refer the reader to (Song et al., 2024b) which provides a general reference on their use for black-box optimization. Relevant to our particular case of regression, works such as (Nguyen and Grover, 2022; Garg et al., 2022) demonstrated the benefits of using raw Transformers as in-context learning (ICL) based regression models, or Neural Processes, with others (Bai et al., 2023; Zhang et al., 2024) established provable guarantees, and (Müller et al., 2022) demonstrated that Transformers trained on synthetic data as “prior-fitted networks” are capable of Bayesian inference, leading to their use in Bayesian optimization (Müller et al., 2023; Nguyen et al., 2023; Nguyen and Grover, 2024).
@@
Unfortunately, as both raw Transformers and GPs require fixed dimensional features, this limits their applications to inputs expressible as tabular features for e.g. hyperparameter tuning, or task-specific embeddings for e.g. chemistry (Maus et al., 2022). Their inherent bias towards smooth functions may also be suboptimal for certain regression tasks (Grinsztajn et al., 2022). Further works have attempted to improve the flexibility of regression-modeling through the use of token-based representations, which allows regressors to be simultaneously used over various types of input formats. Since the context-window of Transformers still remains the most expensive limitation, a useful organization of recent works can be based on their treatment of the total sequence length, roughly equal to:
$\text{(number of trials)}\times\text{(average trial token length)}$ (1)
@@
Chen et al. (2022) use custom tokenizations to minimize trial token length in order to maximize trial count. However, this is restricted to very constrained search spaces (e.g. flat hyperparameter spaces), and lacks flexibility in utilizing arbitrary forms of data. Liu et al. (2024); Vacareanu et al. (2024) use text-to-text chat-based services such as ChatGPT (OpenAI, 2022) and Gemini (Google, 2024) to demonstrate their emergent capabilities for ICL regression, but such methods lack the ability to pretrain over large amounts of offline evaluations, and the use of lengthy natural language can lead to limited trial count. In contrast, Song et al. (2024a); Akhauri et al. (2025) avoid ICL altogether and only places a single trial in the context window to allow arbitrary string representations per trial, but requires the use of alternative but tedious methods of absorbing online data, such as inference-time fine-tuning.
In relation to broad literature such as in LLM reasoning, inference-time compute methods can also be seen through the lens of Bayesian Optimization. For example, verifiers such as Process Reward Models (Lightman et al., 2024) can be seen as ICL-based regression methods over chains of thought, but have only used coarse discretized scores of e.g. $\{-1,0,1\}$ rather than highly precise numeric predictions or uncertainties over vastly different scales.
@@
  • Perform highly precise numeric predictions and uncertainty quantification over diverse objective scales.
This naturally leads to the use of embedding-based methods which can compress any string representation of a candidate into a feature vector, using only a single unit of sequence length when sent to a ICL model such as a raw Transformer. While recent works have applied embeddings for chemical reactions (Kristiadi et al., 2024; Ranković and Schwaller, 2023) and prompt optimization (Hu et al., 2024), no work has assessed their use in the most competitive and widely studied field of standard black-box optimization over tabular-like search spaces, despite evidence (Tang et al., 2025) demonstrating LLM embeddings to possess promising traits such as Lipschitz continuity over tabular features.
@@ 3.1 Preliminaries
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a real-valued function over a search space $\mathcal{X}$. The goal of black-box optimization is to find an input $x^{*}$ which maximizes $f$:
@@
A regressor is a predictive model that estimates a distribution over possible values of $f(\cdot)$ at a given query point $x$, based on the observed history $\{x_{s},y_{s}\}_{s=1}^{t}$ from $t$ previous. Such regressors may also be learnable over additional offline data in addition to the given history.
During inference, the regressor can be transformed into an acquisition function $a:\mathcal{X}\rightarrow\mathbb{R}$ to guide explore-exploit tradeoffs. We assume the existence of a (potentially task-dependent) acquisition optimizer which can quickly and cheaply sample suggestions $x\in\mathcal{X}$, usually in the form of an evolutionary algorithm. The history-dependent acquisition $a_{t+1}(\cdot)$ may then be used to filter out poor candidates or used in an entire Bayesian optimization loop, in which the optimizer identifies the next proposed solution as $x_{t+1}:=\operatorname*{arg\,max}_{x\in\mathcal{X}}a_{t+1}(x)$.
Below in Sections 3.2 and 3.3, we use standard Bayesian optimization modeling techniques as found from previous literature, in particular Transformer Neural Processes (Nguyen and Grover, 2022) and Google Vizier’s (Golovin et al., 2017) Gaussian Process Bandit algorithm (Song et al., 2024c). We emphasize that these specific choices are not the purpose of this work and can be interchangeable with other reasonable regression bodies (Müller et al., 2023) or acquisition functions and optimizers (Balandat et al., 2020b) - the reader is free to skip such details.
@@ 3.2 In-context Transformer Regressor
For our underlying regression model, we use a standard Transformer Neural Process (Nguyen and Grover, 2022), which takes in as input sequence $(\overline{x}_{1}\oplus\overline{y}_{1}),\ldots,(\overline{x}_{t}\oplus\overline{y}_{t})$ where $\overline{y}\in\mathbb{R}^{d}$ is the feature representation of the float $y$ after applying a trainable projection, and $\overline{x}\oplus\overline{y}$ is the trial representation expressed as the concatenation of $\overline{x}$ and $\overline{y}$.
@@
To generate a prediction for a query point $x$, we append the query representation $(\overline{x}\oplus\overline{0})$ to the input sequence where $\overline{0}$ is a dummy value. During the forward pass, the Transformer processes this augmented sequence, allowing the query to attend to all previous trials. The model then produces a parametric output distribution over $\mathbb{R}$ at position $t+1$. Specifically, we assume a Gaussian distribution $\mathcal{N}(\mu_{t+1}(x),\sigma^{2}_{t+1}(x))$, where the mean and standard deviation are predicted by dedicated output heads. Figure 2 provides an overview of the key components of our method.
@@
Parallel Predictions: In order to simultaneously predict over a given set of $k$ target points $x_{t+1},\ldots,x_{t+k}$, we additionally append $(\overline{x}_{t+1}\oplus\overline{0}),\ldots,(\overline{x}_{t+k}\oplus\>\overline{0})$ to the history sequence and generate a custom attention pattern of shape $(t+k)\times t$ where all tokens attend to the history while no tokens attend to the targets. This allows an efficient parallel modeling of $p(y_{t+i}\>|\>x_{t+i},\{x_{s},y_{s}\}_{s=0}^{t})$ for $1\leq i\leq k$ which speeds up training when computing a summed loss over multiple target predictions.
$y$-Normalization: Depending on the function, $y$-value outputs may contain a wide variety of scales, and need to be normalized properly. We use the same normalization procedure as in Google Vizier (Song et al., 2024c), modified to allow incoming target values. These steps consist of, in order: (1) Shifting objectives to have zero mean and divide by standard deviation. (2) Reducing harmful effects of bad outliers by fitting the “bad half” of objectives $\{y_{i}\leq y_{\text{median}}\}$ to a normal curve, using percentiles as z-scores. (3) Linearly scale $y\leftarrow\frac{y-y_{\text{min}}}{y_{\text{max}}-y_{\text{min}}}$ which ensures all historical $y$-values within $[0,1]$ and apply additional damping (e.g. sigmoid or log transform) to target values significantly outside this range.
@@ 3.3 Pretraining and Inference
While the embedder is frozen, we pretrain the weights $\theta$ of the ICL regression Transformer, over all such offline evaluation data. Each training example consists of a sampled task and history cutoff length $t^{\prime}\in[0,T)$ so that $\{x_{s},y_{s}\}_{s\leq t^{\prime}}$ is considered a history, while $\{x_{t^{\prime}+i},y_{t^{\prime}+i}\}_{t^{\prime}+i}^{T}$ are target points, with the loss computed as the mean prediction loss over all targets, i.e.
$\frac{1}{T-t^{\prime}}\sum_{i=1}^{T-t^{\prime}}\ell_{\theta}(x_{t^{\prime}+i},y_{t^{\prime}+i};\{x_{s},y_{s}\}_{s=1}^{t^{\prime}})$ (3)
@@ 3.4 Model Details
In this paper, to demonstrate the validity of our approach on relatively low compute budgets, we intentionally use relatively smaller language model embedder sizes in comparison to the larger and significantly more expensive GPT (OpenAI, 2023) or Gemini (Google, 2024) family of models. Specifically, we use a pretrained T5-XL encoder (1B parameters), based on the encoder-decoder T5-family of models (Raffel et al., 2020). Along with only 8 layers of the underlying regression Transformer, this leads to a maximum required training budget of approximately 16 GPUs for training and 1 GPU for inference, possible with most academic budgets. Further details with respect to model sizes and training hyperparameters can be found in Appendix A.
@@ 4 Experiments
4.1 End-to-End Black-box Optimization
We focus on demonstrating the general applicability of LLM embeddings across a variety of tasks rather than achieving the best possible result against very domain-specific baselines, although we outline possible improvements within specific domains in Section 5. In tabular settings, we purposely use the same algorithmic components (other than replacing the GP regressor) as Google Vizier’s UCB-based “GP-Bandit” (Song et al., 2024c), which allows direct apples-to-apples comparisons. We stress the importance of clean comparison, as it is well known that other components in the Bayesian Optimization pipeline (e.g. choice of acquisition and acquisition optimizer) also strongly affect performance and can lead to confounding factors. GP-Bandit was also reported to be near-optimal across well-known five different industry methods such as Optuna (Akiba et al., 2019) and Ax (Balandat et al., 2020a), and thus achieving competitive results will demonstrate the easy integration of embedding-based regressors into state-of-the-art industry-grade optimization pipelines.
We evaluate the performance of our algorithm on various problems consisting of synthetic, combinatorial, and hyperparameter optimization (exact details in Appendix B). All curves are plotted with mean and 0.5 deviation as error bars over 10 repeats.
@@ 4.1 End-to-End Black-box Optimization
Synthetic Optimization: In common optimization scenarios, the search space is a flat Cartesian product of float and categorical parameter types. Our string-based regression will represent each $x$ with standard JSON over the dictionary mapping parameter names to values, e.g. for a search space with two parameters, one continuous named p0 and another categorical p1, the string representation for an example trial would be {"p0":0.3,"p1":"category_1"}.
We benchmark over the Black-box Optimization Benchmarking (BBOB) suite (ElHara et al., 2019), one of the most widely used synthetic function benchmarks, containing 24 different objectives over continuous search spaces. In order to have a notion of offline “training” data and unseen “test” functions to be optimized, we split the original functions across each landscape type (Separable, Ill-Conditioned, etc.) into training and test sets, and additionally apply randomized transformations (e.g. shifting, rotating, discretizing, increasing/decreasing dimensions) over all objectives to induce non-continuous search spaces with categorical parameters and avoid overfitting. Evaluations used for offline training data were uniformly sampled over the search space.
In Figure 3, we find that Embed-then-Regress is generally comparable with and interestingly can even significantly outperform GP-Bandit in a few cases, such as in Gallagher21Me (7D) and Lunacek (2D). These gaps persist even after 100+ trials, suggesting that the embedding-based regressor can generally lead to a better UCB acquisition function.
@@
Each of these problems can be parameterized by a set of coefficients (e.g. city locations for Travelling Salesman, matrix entries for log-determinant). Note that we are in the bandit setting, in which these coefficients are hidden from the algorithm and the only feedback is the final objective. Similar to before, we thus can also generate offline pretraining data by randomizing these coefficients and problem sizes, and evaluating over random candidates. For our string regression, we may simply use JSON over indices; e.g. {[0]:2, [1]:0, [2]:3, [3]:1} for a permutation space of size 4, e.g. {[0]:1, [1]:3} for a $4\choose 2$} choice space.
@@
Figure 4: ($\uparrow$) Higher is better. Best-so-far curves across 8 randomized combinatorial problems. Title parenthesis $(P)$ means a permutation space of size $P$ and $(N,K)$ denotes a $N\choose K$} choice space. Note that plotting begins at trial 20, since previous trials are random.
@@
While there are few previous works using GPs for e.g. permutation spaces (Deshwal et al., 2022; Oh et al., 2022), they require constructing very domain-specific kernels and complex acquisition optimizers (e.g. semi-definite programming) making them difficult to reproduce. We thus use a simpler optimizer such as Regularized Evolution (Real et al., 2019) which does not need modeling assumptions other than implementing random mutations between trials and can be used broadly (Real et al., 2020). We also empirically found this was better than other evolutionary alternatives such as NSGA-II (Deb et al., 2002) or hill-climbing.
@@
Rather than fully optimizing the acquisition which can be slow for large-scale evolutionary searches, we can simply apply best-of-many sampling by using the regressor’s UCB acquisition to rank sample candidates proposed by evolution, and suggest only the best. In Figure 4, we see that this boosts exploration over the original Regularized Evolution, which can often get stuck at local optima early on.
@@
Figure 5: ($\uparrow$) Higher is better. Best-so-far curves over 8 randomly chosen hyperparameter surrogate functions. Title contains task summary along with number of parameters (#P)(\#P). Normalized objective values are displayed since raw objective values over large ranges, e.g. $y\in[-10^{7},10^{7}]$ from private functions would lack meaningful context.
@@
Figure 6: ($\uparrow$) Higher is better. Violin plots of log-efficiency scores across algorithms over 20 hyperparameter optimization (HPO) surrogate objectives. Embed-then-Regress (EtR) displayed horizontally at Log-Efficiency=0 as reference.
Hyperparameter Optimization: With the advent of industry-wide hyperparameter tuning services (Golovin et al., 2017; Balandat et al., 2020b; Liaw et al., 2018), large amounts of evaluations over expensive but realistic experiments can be stored offline. To efficiently benchmark our method over objectives encountered in the wild, we use surrogate-based benchmarking which has shown to be effective (Zela et al., 2022; Eggensperger et al., 2015) for providing realistic yet cheap comparisons between different algorithms without using huge amounts of compute. Without surrogates, benchmarking could take weeks in wall-clock time and hundreds to thousands of GPU-hours.